At a certain temperature, the pH of pure water is 6.00. Kw for water at this temperature is

1.0 x 10^-6

1.0 x 10^-8

1.0 x 10^-10

1.0 x 10^-12

1.0 x 10^-14
Answer #1

[H3O+] = 10^-pH = 1.0×10^-6 M

Now, pure water contain equal concentration of OH- and H3O+. Thus,

[OH-] = 1.0×10^-6 M

Kw = [H3O+][OH-]

= 1.0×10^-6 × 1.0×10^-6

= 1.0×10^-12 M^2